# Sequential Adiabatic Generation of Chiral Topological States

### v1

@@ Abstract
In previous work, it was shown that non-trivial gapped states can be generated from a product state using a sequential quantum circuit. Explicit circuit constructions were given for a variety of gapped states at exactly solvable fixed points. In this paper, we show that a similar generation procedure can be established for chiral topological states as well, despite the fact that they lack an exactly solvable form. Instead of sequentially applying local unitary gates, we sequentially evolve the Hamiltonian by changing local terms in one subregion and then the next. The Hamiltonian remains gapped throughout the process, giving rise to an adiabatic evolution mapping the ground state from a product state to a chiral topological state. We demonstrate such a sequential adiabatic generation process for free fermion chiral states like the Chern Insulator and the $p+ip$ superconductor. Moreover, we show that coupling a quantum state to a discrete gauge group can be achieved through a sequential quantum circuit, thereby generating interacting chiral topological states from the free fermion ones.
@@ 1 Introduction
In particular, we demonstrate the sequential adiabatic evolution for free fermion chiral states – the Chern insulator in Section 2 and the $p+ip$ superconductor in Section 3 – by showing through an explicit numerical calculation that the Hamiltonian remains gapped in the evolution process. We discuss a scheme based on the coupled wire construction[12, 13], which gives physical intuition for the evolution of the state. The adiabatic process has some similarity to the one used in [14] for the entanglement renormalization transformation of chiral states.
@@ 2.2 Numerical result
Figure 5: The adiabatic evolution of the spectrum as one extra wire is added to the Chern insulator state with width $N_{x}=20$. The spectrum remains gapped as the wire merges into the bulk.
@@ 4 Gauging as a sequential circuit
This circuit is very similar in form to the one discussed in [1] to generate the toric code state from the product state. The only difference is that the Toric Code discussed in [1] is a pure gauge theory (within the low-energy subspace of the Gauss’ law term) while here we include the matter DOF as well.
@@ 5 Sequential adiabatic evolution versus sequential circuits
Some comments are in order about the locality of the sequential adiabatic evolution process. This issue can be discussed either in terms of tuning Hamiltonian couplings, or in terms of the locality properties of a unitary mapping that we apply to the initial product ground state.
@@
Another possible obstruction to removing the tails comes from the fact that it would lead to a state which can be represented by a tensor network with finite bond dimension. It is often stated that such tensor networks states cannot capture chiral states with exponentially decaying correlations in the bulk. This can be proven for free-fermionic tensor networks [17], but recent numerical investigations suggest that they can at least provide faithful approximations [18]. Therefore, the extent to which we can truncate the tails in our unitary evolution while preserving the chiral physics remains an interesting open question.
@@ 6 Discussion
The chiral states covered by methods presented in this paper are either free fermion states or gauge theories with a finite gauge group. The method discussed does not directly apply to states like the chiral Fibbonnaci state. We expect that a similar sequential adiabatic evolution process for such states can be established, although to demonstrate the validity of the process we need to perform strongly correlated numerical simulation.
The sequential quantum circuit for coupling to a finite gauge field works for any symmetric state, gapped, gapless, or even symmetry breaking. It also applies to more general forms of symmetries like higher form symmetry, sub-system symmetry, etc.

### v2

@@ Abstract
In previous work, it was shown that non-trivial gapped states can be generated from a product state using a sequential quantum circuit. Explicit circuit constructions were given for a variety of gapped states at exactly solvable fixed points. In this paper, we show that a similar generation procedure can be established for chiral topological states as well, despite the fact that they lack a zero-correlation-length exactly solvable form. Instead of sequentially applying local unitary gates, we sequentially evolve the Hamiltonian by changing local terms in one subregion and then the next. The Hamiltonian remains gapped throughout the process, giving rise to an adiabatic evolution mapping the ground state from a product state to a chiral topological state. We demonstrate such a sequential adiabatic generation process for free fermion chiral states like the Chern Insulator and the $p+ip$ superconductor. Moreover, we show that coupling a quantum state to a discrete gauge group can be achieved through a sequential quantum circuit, thereby generating interacting chiral topological states from the free fermion ones.
@@ 1 Introduction
In particular, we demonstrate the sequential adiabatic evolution for free fermion chiral states – the Chern insulator in Section 2 and the $p+ip$ superconductor in Section 3 – by showing through an explicit numerical calculation that the Hamiltonian remains gapped in the evolution process. We discuss a scheme based on the coupled wire construction[12, 13], which gives physical intuition for the evolution of the state. The adiabatic process has some similarity to the one used in [14] for the entanglement renormalization transformation of chiral states. In the renormalization procedure of [14], as the length scale gets larger, the procedure involves the adiabatic evolution of nonlocal coupling terms. It is therefore a different setup than the one considered in this paper.
@@ 2.2 Numerical result
Figure 5: The adiabatic evolution of the spectrum as one extra wire is added to the Chern insulator state with width $N_{x}=20$. The spectrum remains gapped as the wire merges into the bulk. In all of the plots, the $x$ axis represents momentum and the $y$ axis represents energy.
@@ 4 Gauging as a sequential circuit
This circuit is very similar in form to the one discussed in [15, 1] to generate the toric code state from the product state. The only difference is that the Toric Code discussed in [15, 1] is a pure gauge theory (within the low-energy subspace of the Gauss’ law term) while here we include the matter DOF as well. Explicitly, if the matter DOFs are originally in the trivial paramagnetic state |$+$|++...$\rangle$| + + … + ⟩, then the final gauge-matter state is exactly the ground state of the $Z_{2}$ lattice gauge theory.
Replacing the operator $X$ in the circuit shown in Fig. 9 with the fermion parity operator $(-1)^{n_{f}}$ gives the fermionic gauging circuit that couples a fermionic state to the $Z_{2}$ gauge field. In particular, in the contolled-Not gate (controlled by a $\tau_{x}$), the $X$ operator is replaced with the fermionic parity operator $(-1)^{n_{f}}$. The fermion parity operator is fermion even. Therefore, the replacement does not lead to any issue of non-commutativity among operators.
@@ 5 Sequential adiabatic evolution versus sequential circuits
Some comments are in order about the locality of the sequential adiabatic evolution process. Consider generating the chiral state on a flat 2D plane. This issue can be discussed either in terms of tuning Hamiltonian couplings, or in terms of the locality properties of a unitary mapping that we apply to the initial product ground state.
@@
Note that exponentially decaying correlation alone does not forbid the possibility of sequential circuit generation. Indeed, it is known that matrix product states can all be generated using sequential circuits, even if they have nonzero correlation length[2, 3]. Another possible obstruction to removing the tails in the chiral state generating process comes from the fact that it would lead to a state which can be represented by a tensor network with finite bond dimension. It is often stated that such tensor networks states cannot capture chiral states with exponentially decaying correlations in the bulk. This can be proven for free-fermionic tensor networks [18], but recent numerical investigations suggest that they can at least provide faithful approximations [19]. Therefore, the extent to which we can truncate the tails in our unitary evolution while preserving the chiral physics remains an interesting open question.
@@ 6 Discussion
The chiral states covered by methods presented in this paper are either free fermion states or gauge theories with a finite gauge group. The method discussed does not directly apply to states like the chiral Fibonacci state. We expect that a similar sequential adiabatic evolution process for such states can be established, although to demonstrate the validity of the process we need to perform strongly correlated numerical simulation.
The sequential quantum circuit for coupling to a finite gauge field works for any symmetric state, gapped, gapless, or even symmetry breaking. It also applies to more general forms of symmetries like higher form symmetry, sub-system symmetry, etc. Therefore, this kind of circuit should be useful for further investigations of the gauging of many-body systems, including its implementation on quantum devices.